OpenStudy (fibonaccichick666):

Graph Theory assistance please!! A connected graph G contains k blocks and l cut vertices. What is the relationship between k and l. So I think it should be that k=l+1 But I am not sure if that is right or how to prove it.

OpenStudy (fibonaccichick666):

it might be k=l+2

OpenStudy (kainui):

I don't really know what the words mean exactly, but if I have one loaf of bread, it has zero cuts and if I slice it in half I have two pieces and one cut. So pieces = cuts + 1 seems right to me based on this thinking haha.

OpenStudy (fibonaccichick666):

it definitely is not k=l+1 I now ewalize

OpenStudy (fibonaccichick666):

right thinking, but there are two ends to every piece of bread right?

OpenStudy (fibonaccichick666):

I just found a theorem, and a counter example

OpenStudy (fibonaccichick666):

that is the only reason I now know

OpenStudy (kainui):

I mean I am not entirely sure what we're talking about here, it's just my guess haha. I don't know what blocks or cut vertices are.

OpenStudy (fibonaccichick666):

oh, yea, that makes things more difficult haha

OpenStudy (fibonaccichick666):

so here listen to this "Every two distinct blocks of G have at most one vertex in common; and if they have a vertex in common, then this vertex is a cut-vertex of G"

OpenStudy (kainui):

What's a distinct block of G?

OpenStudy (fibonaccichick666):

on a graph with 3 cut vertices, we had 5 blocks

OpenStudy (fibonaccichick666):

a block is the biggest section you can cut off from a graph at the cut vertex

OpenStudy (kainui):

So basically that means that these are like the final strand left to separate the graph into two separate graphs then?

OpenStudy (fibonaccichick666):

eg

OpenStudy (fibonaccichick666):

so u,w and y are cut vertices

OpenStudy (kainui):

I think I see now.

OpenStudy (kainui):

So the number of edges at the cut vertices will be the number of graphs we get after we cut it there I think.

OpenStudy (fibonaccichick666):

the blocks are